# Query-Specific GNN: A Comprehensive Graph Representation Learning Method for Retrieval Augmented Generation

### v1

@@ 3.2 Query-Specific Graph Neural Network
$\alpha_{i,j}=Sim\left(\mathbf{h}^{l-1}_{i}W^{q}_{\alpha},\mathbf{h}^{l-1}_{j}W^{k}_{\alpha}\right),$ (2)
@@
$attn=Softmax\left(\left\{\alpha_{i,j}+\beta_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
$msg_{i}=\sum\limits_{j\in\mathcal{N}(i)}attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
@@
where $\mathbf{h}^{l}_{i/j}\in R^{n}$ is the representation of node $i/j$ in layer $l$, initialized with text embeddings from an embedding model  [^3] Embeddings are compressed to dimension $n$ via linear transformations as information bottlenecks.. $Sim(.)$ calculates the cosine similarity. $\mathbf{q}\in R^{n}$ is the query embedding.||$|$ represents the vector concatenation. $W^{q}_{\alpha/\beta}\in R^{n\times n}$, $W^{k}_{\alpha}\in R^{n\times n}$, $W^{k}_{\beta}\in R^{2n\times n}$, $W^{v}\in R^{n\times n}$ are all learnable parameters. $Norm(.)$ is the normalization function, $P(.)$ is the 2-layers MLP with ReLU  (Glorot et al., 2011).
Intra-level message passing captures two types of relationships in the Multi-L KG. One is the basic semantic relationship encoded in $\mathcal{E}_{oo}$ and the other is the logical coherence relationship represented by $\mathcal{E}_{cc}$. The aggregation considers not only connectivity, but also semantic similarity $\alpha_{i,j}$ and query alignment $\beta_{i,j}$, which minimizes noise while ensuring query-aware aggregation.
@@
$\mathbf{p}_{i,j}=\left((\mathbf{h}^{l-1}_{i})W^{t}||(\mathbf{h}^{l-1}_{j})W^{s}\right),$ (3)
@@
$attn=Softmax\left(\left\{\gamma_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
$\mathbf{msg}_{i}=\sum\limits_{j\in\mathcal{N}(i)}attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
@@
The inter-level message passing fuses the local, global relationships from hierarchical levels into representations, which contains comprehensive understanding towards the query.
Based on these two processes, QSGNN calculates the representations as follows:
$\mathbf{H}^{l}_{intra}=IntraMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{E}_{oo},\mathcal{E}_{cc},\mathbf{H}^{l-1}),$ (4)
$\mathbf{H}^{l}=InterMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{C},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cd},\mathbf{H}^{l}_{intra}),$
@@
3.3 Training Strategy for QSGNN
To enhance the representation learning ability of QSGNN, we first pre-train it on synthesized data and then we fine-tune it with human annotations. The pre-training data are the synthesized (question, document) pairs extracted from the same corpora used to construct the Multi-L KG. These pairs contain one-hop questions and two-hop questions, which are all generated by OpenIE  [^4] The prompts can be found in Appendix  A.11.
One-hop Question. We generate one-hop questions based on the triples extracted from documents. For each triple-document pair $\{(sbj,verb,obj),d\}$, the question is generated in form of $(?,verb,obj)$ or $(sbj,verb,?)$ where the answer is sbjsbj or objobj and the support document is $d$.
Two-hop Question. Two-hop questions are generated based on relation chains formed by shared entities across different documents. For instance, if document $d_{i}$ contains the triple $(sbj_{i},verb_{i},ent)$ and document $d_{j}$ contains the triple $(ent,verb_{j},obj_{j})$ where entent is the common entity, we form the relation chain as $(sbj_{i},verb_{i},ent,verb_{j},obj_{j})$, the corresponding question is generated in the form $(?,verb_{i},ent,verb_{j},obj_{j})$ or $(sub_{i},verb_{i},ent,verb_{j},?)$ where the answer is $sub_{i}$ or $obj_{j}$ and the support documents are $d_{i}$ and $d_{j}$.
The QSGNN is pre-trained on the synthesized data using the NT-Xent loss  (Chen et al., 2020b):
$\mathcal{L}_{\text{NT-Xent}}=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{i})/\tau)}{\sum_{j\in Neg(i)}\mathbf{1}_{[j\neq i]}\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{j})/\tau)},$ (5)
where $\mathbf{q}_{i}$ denotes the query embedding, $h_{i}$ is the representation of support document, $M$ is the batch size, $\tau$ is the temperature parameter and we set it as 1.0 in our implementation, $Neg(.)$ denotes the negative sampling and we employ a hard negative sampling strategy  (Schroff et al., 2015; Xu et al., 2022) with sampling number set as 30.
After pre-training, we fine-tune QSGNN on the downstream retrieval task. In this stage, the questions are collected by humans with answers and support documents. We fine-tune QSGNN with the same loss function as Equation  5.
3.4 QSGNN Retrieval and Generation
Given a query $\mathbf{q}$, the retrieval process of QSGNN is denoted as:
$score=Sim(\mathbf{q},\mathbf{H}^{l}),$ (6)
@@ 3.4 QSGNN Retrieval and Generation
$response=LLM(query,\mathcal{R}).$ (7)
4 Experimental Results
4.1 Experimental Settings
Dataset. We evaluate the effectiveness of our proposed method on three multi-hop QA benchmarks: MuSiQue  (Trivedi et al., 2022a), 2WikiMultiHopQA(2Wiki)  (Ho et al., 2020) and HotpotQA  (Yang et al., 2018). For a fair comparison, we use the same evaluation set(1000 samples for each dataset) as prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025). Additionally, we sample extra 1,000 questions from the original dataset for fine-tuning and 225 questions for testing, without overlap with the evaluation set. The basic statistics of these dataset are shown in Table  1. Details about data process are presented in Appendix  A.2.
Table 1: Dataset statistics.
@@ 4.1 Experimental Settings
Baselines. We compare our method (QSGNN) with four baseline categories: i) Naive Retriever: we use BM25 (Robertson & Walker, 1994) and Contriever (Izacard et al., 2021) together with IRCoT (Trivedi et al., 2022b) as the naive retrievers. ii) Text Embedding: we use NV-Embed-v2-7B  (Lee et al., 2024) and GTE-Qwen2-7B-Instruct (Li et al., 2023) to generate embeddings for retrieval. iii) Graph Search: we use GraphRAG  (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), LightRAG (Guo et al., 2024) HippoRAG (Jimenez Gutierrez et al., 2024) and HippoRAG2 (Gutiérrez et al., 2025) as the graph search baselines. iv) GNN Retriever: we use GNN-RAG (Mavromatis & Karypis, 2024) and GFM-RAG  (Luo et al., 2025) as the GNN retriever baselines. The implementation details for all baselines can be found in Appendix  A.3.
Metrics. Follow prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025), we evaluate retrieval performance using recall@2 and recall@5 (top-5 retrieved documents) and Question-Answering(QA) performance with exact match (EM) and F1 scores.
Implementation Details. Our QSGNN is implemented with 2 layers, where each layer has one intra-leverl and one inter-level message passing process. Information dimension is set as 128. All the training uses 2 NVIDIA A100(80G) GPUs. In the pre-training stage, we use 95% samples for training and 5% samples for checkpoint selection [^5] The checkpoint is selected based on the retrieval performance on the 5% samples using Recall@5.. The max epoch number is set as 5. We set learning rate as 1e-4 without weight decay. The checkpoint is saved every 2000 steps. In the fine-tuning stage, the max epoch number is set as 3. We set learning rate as 5e-4. The checkpoint is saved every 100 steps and we select checkpoint based on test set (225 samples) to evaluate the performance. We use NV-Embed-v2-7B  (Lee et al., 2024) as the embedding model and we use Llama-3.3-70B-Instruct  (AI@Meta., 2024) as the OpenIE model. We retrieve the top-5 documents for LLM(Llama-3.3-70B-Instruct and GPT-4o-mini  (OpenAI., 2024)) as context for QA task.
4.2 Results and Discussion
Retrieval Task. Table 2 summarizes the retrieval performance across baselines. We have following key observations: i) QSGNN outperforms all baseline methods, especially compared to other GNN based methods, it achieves +11.9%, +12.9% and +5.4% improvements on MuSiQue, 2Wiki and HotpotQA respectively and +9.8% average improvement, showing its effectiveness on multi-hop questions. ii) The classic retrieval algorithms (BM25/Contriever) and text embedding baselines have limited performance as they fail to capture document relationships which is crucial for answering multi-hop questions. iii) Although GNN based methods are designed to capture multi-hop relationships, their message passing mechanisms may inadvertently aggregate irrelevant information, which degrades retrieval effectiveness. iv) Graph search methods generally perform better than other baselines, but their performance is still constrained by inherent limitations. The heuristic exploration rules often overlook multi-level information and may favor specific path patterns, resulting in data bias. For instance, the performance of RAPTOR varies a lot since it leverages Gaussian Mixture Model to detect clusters in KGs, which is highly biased on data.
@@ 4.2 Results and Discussion
Table 2: Retrieval performance of the baselines. Rec@2/Rec@5 denotes Recall@2/Recall@5. Average means the average performance across all the datasets. We highlight the best results with bold and the second best results with under line. We do not report the result of LightRAG or GraphRAG because they do not retrieve documents.
MuSiQue 2Wiki HotpotQA Average
Method Rec@2 Rec@5 Rec@2 Rec@5 Rec@2 Rec@5 Rec@2 Rec@5
BM25 32.76 43.62 55.32 65.31 57.36 75.82 48.48 61.58
Contriever+IRCoT 34.92 46.27 46.6 57.53 56.21 74.37 45.91 59.39
GTE-Qwen2-7B 51.24 67.72 66.73 77.78 77.24 90.15 65.07 78.55
NV-Embed-v2 50.12 68.19 67.16 78.66 79.27 92.82 65.52 79.89
GNN-RAG 34.13 53.72 58.67 69.14 67.18 81.69 53.33 68.18
GFM-RAG 49.66 67.19 62.92 79.20 74.05 88.86 62.21 78.42
RAPTOR 47.68 69.55 66.45 82.67 77.83 93.19 63.99 81.80
HippoRAG 47.55 68.15 71.95 86.25 70.17 84.11 63.22 79.50
HippoRAG2 50.42 72.29 73.31 88.91 78.25 92.56 67.33 84.59
QSGNN(Ours) 53.88 75.23 74.02 89.47 78.42 93.67 68.77 86.12
Table 3: QA performance using Llama-3.3-70B-Instruct. No Retrieval means the LLM response without external knowledge. Average means the average performance across all the datasets. We highlight the best results with bold and the second best results with under line.
MuSiQue 2Wiki HotpotQA Average
Method EM F1 EM F1 EM F1 EM F1
No Retrieval 17.29 26.26 36.37 41.87 36.42 46.69 30.03 38.27
BM25 20.38 28.04 43.15 49.49 51.26 61.44 38.26 46.32
Contriever+IRCoT 24.29 31.62 38.02 40.90 49.89 60.30 37.4 44.27
GTE-Qwen2-7B 33.60 40.9 54.11 59.61 58.61 70.97 48.77 57.16
NV-Embed-v2 31.66 39.26 53.78 59.99 59.53 72.27 48.32 57.17
GNN-RAG 28.29 34.09 38.46 49.77 52.03 63.52 39.59 49.13
GFM-RAG 30.05 39.06 48.99 59.58 55.60 67.86 44.88 55.5
RAPTOR 32.46 40.30 48.57 62.15 60.21 73.25 47.08 58.57
LightRAG 6.18 13.66 6.75 22.50 12.82 26.25 8.58 20.80
GraphRAG 34.10 41.90 52.46 63.63 59.89 73.35 48.82 59.63
HippoRAG2 35.16 42.44 56.60 66.18 60.92 74.06 50.89 60.89
QSGNN(Ours) 36.65 44.93 57.02 66.83 60.23 74.44 51.3 62.07
QA Task. We conduct QA experiments across the methods and the results are shown in Table 3. QSGNN outperforms other baselines due to its accurate retrieval. No Retrieval achieves low EM and F1 scores, demonstrating the importance of external knowledge for the LLM’s response. Notably, LightRAG struggles to perform well on these datasets, which is consistent with other reports  (Gutiérrez et al., 2025; Luo et al., 2025). The main reason is that LightRAG is incompatible with multi-hop retrieval workflows as discussed in  (Luo et al., 2025). The performance of the other methods have the consistent observations with Table 2 because the retrieval accuracy directly determines the quality of response. The results of GPT-4o-mini are shown in Appendix  A.4.
Multi-hop Performance. In order to analyze the detailed performance on multi-hop questions, we evaluate the performance of the competitive baselines on different hop numbers. The results of MuSiQue are shown in Table 4 and Figure  1. Firstly, we find that all the methods exhibit performance decline as hop number increases, because of the exponentially growing search space and noise accumulation. Secondly, GNN based methods particularly underperform in high-hop scenarios due to: i) the message passing aggregates the information of irrelevant nodes, ii) they fail to consider multiple information levels. Finally, as for the QSGNN, it achieves competitive performance in 2-hop questions compared to other baselines and the improvements become noticeable as the hop number increases, where it gets +33.8%, +85.8% and +231% improvements in Recall@5, F1 and EM on 4-hop questions respectively. This advantage stems from QSGNN’s query-guided attention mechanism and multi-level message passing, which significantly reduce the impact of noise and fuse the multi-information levels comprehensively(see the ablation study). The result of 2Wiki can be found in Appendix  A.5.

### v2

@@ 3.2. Question-Adaptive Graph Neural Network
(2) $\alpha_{i,j}=Sim\left(\mathbf{h}^{l-1}_{i}W^{q}_{\alpha},\mathbf{h}^{l-1}_{j}W^{k}_{\alpha}\right),$
@@
$intra\_attn=Softmax\left(\left\{\alpha_{i,j}+\beta_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
$msg_{i}=\sum\limits_{j\in\mathcal{N}(i)}intra\_attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
@@
where $\mathbf{h}^{l}_{i/j}\in R^{n}$ is the representation of node $i/j$ in layer $l$, initialized with text embeddings from an embedding model  [^3] Embeddings are compressed to dimension $n$ via linear transformations as information bottlenecks.. $Sim(.)$ calculates the cosine similarity. $\mathbf{q}\in R^{n}$ is the question embedding.||$|$ represents the vector concatenation. $W^{q}_{\alpha/\beta}\in R^{n\times n}$, $W^{k}_{\alpha}\in R^{n\times n}$, $W^{k}_{\beta}\in R^{2n\times n}$, $W^{v}\in R^{n\times n}$ are all learnable parameters. $Norm(.)$ is the normalization function, $P(.)$ is the 2-layers MLP with ReLU  (Glorot et al., 2011).
Intra-level message passing captures two types of relationships in the Multi-L KG. One is the basic semantic relationship encoded in $\mathcal{E}_{oo}$ and the other is the logical coherence relationship represented by $\mathcal{E}_{cc}$. The aggregation considers not only connectivity, but also semantic similarity $\alpha_{i,j}$ and question alignment $\beta_{i,j}$, which minimizes noise while ensuring question-aware aggregation.
@@
(3) $\mathbf{p}_{i,j}=\left((\mathbf{h}^{l-1}_{i})W^{t}||(\mathbf{h}^{l-1}_{j})W^{s}\right),$
@@
$inter\_attn=Softmax\left(\left\{\gamma_{i,j}|j\in\mathcal{N}(i)\right\}\right),$
$\mathbf{msg}_{i}=\sum\limits_{j\in\mathcal{N}(i)}inter\_attn_{i,j}\mathbf{h}^{l-1}_{j}W^{v},$
@@
The inter-level message passing fuses the local, global relationships from hierarchical levels into representations, which contains comprehensive understanding towards the question.
Based on these two processes, Quest-GNN calculates the representations as follows:
(4) $\mathbf{H}^{l}_{intra}=IntraMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{E}_{oo},\mathcal{E}_{cc},\mathbf{H}^{l-1}),$
$\mathbf{H}^{l}=InterMQ(\mathbf{q},\mathcal{O},\mathcal{C},\mathcal{D},\mathcal{E}_{oc},\mathcal{E}_{od},\mathcal{E}_{cd},\mathbf{H}^{l}_{intra}),$
@@
3.3. Training Strategy for Quest-GNN
To enhance the representation learning ability of Quest-GNN, we first pre-train it on synthesized data and then we fine-tune it with human annotations. The pre-training data are the synthesized (question, document) pairs extracted from the same corpora used to construct the Multi-L KG. These pairs contain one-hop questions and two-hop questions, which are all generated by OpenIE  [^4] The prompts can be found in Appendix  A.5.
One-hop Question. We generate one-hop questions based on the triples extracted from documents. For each triple-document pair $\{(sbj,verb,obj),d\}$, the question is generated in form of $(?,verb,obj)$ or $(sbj,verb,?)$ where the answer is sbjsbj or objobj, the support document is $d$ and the ?? denotes the question target asking the relation with $sbj/obj$.
Two-hop Question. Two-hop questions are generated based on relation chains formed by shared entities across different documents. For instance, if document $d_{i}$ contains the triple $(sbj_{i},verb_{i},ent)$ and document $d_{j}$ contains the triple $(ent,verb_{j},obj_{j})$ where entent is the common entity, we form the chain as $(sbj_{i},verb_{i},ent,verb_{j},obj_{j})$, the corresponding question is generated from the relation chain in the form $(?,verb_{i},ent,verb_{j},obj_{j})$ or $(sbj_{i},verb_{i},ent,verb_{j},?)$ where the question is about the two-hop relationship between $sbj_{i}$ and $obj_{j}$, the support documents are $d_{i}$ and $d_{j}$ and the answer is $sbj_{i}$ or $obj_{j}$.
The Quest-GNN is pre-trained on the synthesized data using the NT-Xent loss  (Chen et al., 2020b):
(5) $\mathcal{L}_{\text{NT-Xent}}=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{i})/\tau)}{\sum_{j\in Neg(i)}\mathbf{1}_{[j\neq i]}\exp(\text{sim}(\mathbf{q}_{i},\mathbf{h}_{j})/\tau)},$
where $\mathbf{q}_{i}$ denotes the question embedding, $h_{i}$ is the representation of support document, $M$ is the batch size, $\tau$ is the temperature parameter and we set it as 1.0 in our implementation, $Neg(.)$ denotes the negative sampling and we employ a hard negative sampling strategy  (Schroff et al., 2015; Xu et al., 2022) with sampling number set as 30.
After pre-training, we fine-tune Quest-GNN on the downstream retrieval task. In this stage, the questions are collected by humans with answers and support documents. We fine-tune Quest-GNN with the same loss function as Equation  5.
3.4. Quest-GNN Retrieval and Generation
Given a question $\mathbf{q}$, the retrieval process of Quest-GNN is denoted as:
(6) $score=Sim(\mathbf{q},\mathbf{H}^{l}),$
@@ 3.4. Quest-GNN Retrieval and Generation
(7) $response=LLM(question,\mathcal{R}).$
4. Experimental Results
4.1. Experimental Settings
4.1.1. Dataset
We evaluate the effectiveness of our proposed method on three multi-hop QA benchmarks: MuSiQue  (Trivedi et al., 2022a), 2WikiMultiHopQA(2Wiki)  (Ho et al., 2020) and HotpotQA  (Yang et al., 2018). For a fair comparison, we use the same evaluation set (1000 samples for each dataset) as prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025). Additionally, we sample extra 1,000 questions from the original dataset for fine-tuning and 225 questions for validation, without overlap with the evaluation set. The basic statistics of these dataset are shown in Table  1 and detailed statistics and sampling strategy are presented in Appendix  A.1.
Table 1. Dataset statistics.
@@ 4.1.1. Dataset
Table 2. Retrieval and QA performance using Llama-3.3-70B-Instruct. Rec@2/Rec@5 denotes Recall@2/Recall@5. No Retrieval means the LLM response without external knowledge. We highlight the best results with bold and the second best results with under line. We do not report the recall of LightRAG or GraphRAG because they do not retrieve documents.
MuSiQue 2Wiki HotpotQA
Method Rec@2 Rec@5 EM F1 Rec@2 Rec@5 EM F1 Rec@2 Rec@5 EM F1
No Retrieval - - 17.29 26.26 - - 36.37 41.87 - - 36.42 46.69
BM25 32.76 43.62 20.38 28.04 55.32 65.31 43.15 49.49 57.36 75.82 51.26 61.44
Contriever+IRCoT 34.92 46.27 24.29 31.62 46.6 57.53 38.02 40.90 56.21 74.37 49.89 60.30
GTE-Qwen2-7B 51.24 67.72 33.60 40.9 66.73 77.78 54.11 59.61 77.24 90.15 58.61 70.97
NV-Embed-v2 50.12 68.19 31.66 39.26 67.16 78.66 53.78 59.99 79.27 92.82 59.53 72.27
Auto-RAG 41.37 61.40 27.36 34.89 55.63 74.94 45.28 55.41 65.81 79.50 51.21 60.33
Self-RAG 38.42 59.17 25.83 34.75 52.92 72.45 44.76 53.27 66.36 79.27 51.45 60.02
KiRAG 42.96 63.05 28.66 36.35 58.67 79.01 48.91 59.53 67.18 81.69 52.03 63.52
GNN-RAG 34.13 53.72 28.29 34.09 58.67 69.14 38.46 49.77 67.18 81.69 52.03 63.52
G-retriever 47.14 66.46 30.13 38.04 62.38 78.05 46.84 58.10 74.16 89.49 54.73 65.90
GFM-RAG 49.66 67.19 30.05 39.06 62.92 79.20 48.99 59.58 74.05 88.86 55.60 67.86
GraphRAG - - 34.10 41.90 - - 52.46 63.63 - - 59.89 73.35
LightRAG - - 6.18 13.66 - - 6.75 22.50 - - 12.82 26.25
RAPTOR 47.68 69.55 32.46 40.30 66.45 82.67 48.57 62.15 77.83 93.19 60.21 73.25
ToG2 46.16 68.04 32.75 40.28 64.86 81.22 47.74 62.14 75.14 90.18 57.99 72.53
HippoRAG2 50.42 72.29 35.16 42.44 73.31 88.91 56.60 66.18 78.25 92.56 60.92 74.06
Quest-GNN(Ours) 53.88 75.23 36.65 44.93 74.02 89.47 57.02 66.83 78.42 93.67 60.23 74.44
4.1.2. Baselines
We compare our method (Quest-GNN) with five baseline categories:
  • Naive Retriever: we use BM25 (Robertson and Walker, 1994) and Contriever (Izacard et al., 2021) together with IRCoT (Trivedi et al., 2022b) as the naive retrievers.
  • Text Embedding: we use NV-Embed-v2-7B  (Lee et al., 2024) and GTE-Qwen2-7B-Instruct (Li et al., 2023) to generate embeddings for retrieval.
  • Iterative Retriever: we use Self-RAG  (Asai et al., 2024), Auto-RAG (Yu et al., 2024) and KiRAG  (Fang et al., 2025) as the multi-step RAG baselines.
  • Graph Search: we use GraphRAG  (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), LightRAG (Guo et al., 2024), ToG2 (Ma et al., 2024) and HippoRAG2 (Gutiérrez et al., 2025) as the graph search baselines.
  • GNN Retriever: we use GNN-RAG (Mavromatis and Karypis, 2024), G-retriever (He et al., 2024) and GFM-RAG  (Luo et al., 2025) as the GNN retriever baselines.
The settings of baselines can be found in Appendix  A.2.
4.1.3. Metrics
Follow prior works  (Gutiérrez et al., 2025; Jimenez Gutierrez et al., 2024; Luo et al., 2025), we evaluate retrieval performance using recall@2 and recall@5 (top-5 retrieved documents) and Question-Answering(QA) performance with exact match (EM) and F1 scores.
4.1.4. Implementation Details
Our Quest-GNN is implemented with 2 layers, where each layer has one intra-leverl and one inter-level message passing process. Information dimension is set as 128. All the training uses 2 NVIDIA A100(80G) GPUs. In the pre-training stage, we use 95% samples for training and 5% samples for checkpoint selection [^5] The checkpoint is selected based on the retrieval on the 5% samples using Recall@5.. The max epoch number is set as 5. We set learning rate as 1e-4 without weight decay. The checkpoint is saved every 2000 steps. In the fine-tuning stage, the max epoch number is set as 3. We set learning rate as 5e-4. The checkpoint is saved every 100 steps and we select checkpoint based on dev set (225 samples) to evaluate the performance. We use NV-Embed-v2-7B  (Lee et al., 2024) as the embedding model and we use Llama-3.3-70B-Instruct  (AI@Meta., 2024) as the OpenIE model. We retrieve the top-5 documents for LLM(Llama-3.3-70B-Instruct and GPT-4o-mini  (OpenAI., 2024)) as context for QA task.